brian and ryan have ages in the ratio 2:3. in 4 years time the ratio will be 5:7. what are their current ages?


Answer by hk(4)

let's assume age of brian and ryan as x'and 'y'resply.from information given,
x/y=2/3 or y=1.5 x
in 4 years, (x+4)/(y+4) will be 5/7. Solving for x, we get
(x+4)/(1.5x+4)=5/7 or 7x+28= 7.5x + 20 i.e. x=16. Hence y is 24.